• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 07:42:57 on November 7,2000

    Entry number 51650

    This entry is a followup to: 51646

    keyword=Beam and DAQ Problems Fixed!

    The problems with the beam were fixed and beam was restored to the Hall at 07:00.
    By 07:30 Bob helped lead us to the problem in the right arm, which turned out to
    be the HV had switched off. The voltage was restored and we resumed taking data.
    The problem with the FPP portserver has occured before and Bob will replace the
    unit during the machine recovery today.